OpenStudy (gabylovesyou):

@jim_thompson5910 how do i simplify this equation to find center and radius x2 − 4x + y2 + 8y = −4 The center is located at (−2, −4), and the radius is 4. The center is located at (2, −4), and the radius is 4. The center is located at (−2, −4), and the radius is 16. The center is located at (2, −4), and the radius is 16.

OpenStudy (anonymous):

Do you know "Completion Of Square Method" ??

OpenStudy (gabylovesyou):

no

OpenStudy (anonymous):

The equation of a circle on a coordinate plane is \[(x-h)^{2}+(y-k)^{2}=r ^{2}\]

OpenStudy (anonymous):

h and k are the x and y coordinates of the center of the circle

OpenStudy (anonymous):

@arilove1d she does not know completion of square method..

OpenStudy (anonymous):

Do you want to learn the method @Gabylovesyou ??

OpenStudy (anonymous):

\[x^2+y^2-4x+8y+4=0\] compare with \[x^2+y^2+2gx+2fy+c=0\] center is (-g,-f) and radius is \[\sqrt{g^2+f^2-c}\]

OpenStudy (gabylovesyou):

@waterineyes sure

OpenStudy (anonymous):

or you can use completing square method as suggested above.

OpenStudy (anonymous):

@surjithayer great, I was earlier thinking the same when she said, she does not know the method.. :)

OpenStudy (anonymous):

@Gabylovesyou can you use surjit's method? Can you apply?

OpenStudy (gabylovesyou):

im looking at it

OpenStudy (anonymous):

Take your time. we are in no hurry.. :)

OpenStudy (gabylovesyou):

yea i have no idea.. i was trying to do my lessons example.... Step 1 [original equation]: x^2 - 4x + y^2 + 8y = -4 Step 2 [group like terms]: (x^2 - 4x) + (y^2 + 8y) = -4 Step 3 [complete the quadratics]:(x^2 - 4x + 4) + (y2 + 8y + 16) = -4 + (4 + 8)

OpenStudy (gabylovesyou):

idk if i did step 3 correctly...

OpenStudy (anonymous):

What is this?

OpenStudy (gabylovesyou):

this is the way my lesson teaches it

OpenStudy (anonymous):

Is this not Completion of square method?

OpenStudy (gabylovesyou):

i think it is

OpenStudy (gabylovesyou):

yes .. i just dont know how to do it lol

OpenStudy (anonymous):

Step 3 [complete the quadratics]: You yourself typed it.. :P

OpenStudy (phi):

complete the quadratics means the same as complete the square

OpenStudy (gabylovesyou):

oh.. did i do it right ?

OpenStudy (anonymous):

This is as simple as you log in to OS daily... :P

OpenStudy (anonymous):

Okay, just one mistake.. You must add 16 on right hand side to balance your equation.. :)

OpenStudy (gabylovesyou):

didnt i do that ?

OpenStudy (anonymous):

Step 3 [complete the quadratics]:(x^2 - 4x + 4) + (y2 + 8y + 16) = -4 + (4 + 8) Not 8 in last but 16.. Step 3 [complete the quadratics]:(x^2 - 4x + 4) + (y2 + 8y + 16) = -4 + (4 + 16)

OpenStudy (anonymous):

You did (4+8), you should do (4+16), getting?

OpenStudy (gabylovesyou):

ohh ok .. then it says this Step 4 [simplify the equation]: (x2 − 4x+4) + (y2 + 8y+16) = 16 Step 5 [factor each quadratic]: (x - 2)^2 + (y + 4)^2 = 16^2 Step 6 [identify the center and radius]: (2,-4) radius 16 ?

OpenStudy (gabylovesyou):

so answer is D?

OpenStudy (anonymous):

Do you still think radius is 16 ??

OpenStudy (anonymous):

How you changed \(16\) to \(16^2\) there?

OpenStudy (gabylovesyou):

omg i ment 4^2 cx

OpenStudy (anonymous):

That is right, don't do silly mistakes, go slow but correct.. :)

OpenStudy (gabylovesyou):

so answer is B....

OpenStudy (anonymous):

So, \(B\) ??

OpenStudy (anonymous):

yep.. :) Good.. :)

OpenStudy (gabylovesyou):

yes... lol thanks ! ;p